OpenStudy (anonymous):

What is the solution to the equation 3^(2x) ≈ 4 ?

OpenStudy (anonymous):

9^x = 4 ln(9^x) = ln(4) xln(9) = ln(4) x = ln(4)/ln(9)

OpenStudy (anonymous):

would it be 0.631?

OpenStudy (uri):

Honey,let me check.

OpenStudy (uri):

right.

OpenStudy (anonymous):

thanks